# MixGRPO: Unlocking Flow-based GRPO Efficiency with Mixed ODE-SDE

### v2

@@ 3.1 Mixed ODE-SDE Sampling in GRPO
In the RF framework, the model is used to predict the velocity field of the deterministic ODE, represented as $\mathbf{v}_{\theta}(\mathbf{x}_{t},t)=\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}$. Following Liu et al. (2025a), the score function is represented as $\nabla_{\mathbf{x}_{t}}\log{q_{t}(\mathbf{x}_{t})}=-\frac{\mathbf{x}_{t}}{t}-\frac{1-t}{t}\mathbf{v}_{\theta}(\mathbf{x}_{t},t)$. The $g(t)$ is represented as the standard deviation of the noise $g(t)=\sigma_{t}$. According to the definition of the standard Wiener process, we use $\mathrm{d}\mathbf{w}=\sqrt{\mathrm{d}t}\epsilon$, where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$. Applying Euler-Maruyama discretization for SDE and Euler discretization for ODE, we build the final denoising process in MixGRPO:
[…]
According to Eq. (1), we only need to optimize GRPO (Shao et al., 2024) at the time within the interval $S$ for $N$ samples in the group. The final training objective is represented as follows:
$\mathcal{J}_{\text{MixGRPO}}(\theta)$ $=\mathbb{E}_{c\sim\mathcal{C},~\{\mathbf{x}^{i}_{T}\}^{N}_{i=0}\sim\pi_{\theta_{\text{old}}}(\cdot|c)}$ (8)
$\left[\frac{1}{N}\sum^{N}_{i=1}\frac{1}{|S|}\sum_{t\in S}\text{min}\left(r^{i}_{t}(\theta)A^{i},\text{clip}\left(r^{i}_{t}(\theta),1-\varepsilon,1+\varepsilon\right)A^{i}\right)\right],$
where $r^{i}_{t}(\theta)$ is referred to as the policy ratio and $A^{i}$ is the advantage score. We set $\varepsilon=0.0001$,
$r^{i}_{t}(\theta)=\frac{q_{\theta}(\mathbf{x}_{t+\Delta t}|\mathbf{x}_{t},c)}{q_{\theta_{\text{old}}}({\mathbf{x}_{t+\Delta t}|\mathbf{x}_{t},c})}\quad\text{and}\quad A^{i}=\frac{R\left(\mathbf{x}^{i}_{T},c\right)-\text{mean}\left(\{R\left(\mathbf{x}^{i}_{T},c\right)\}^{N}_{i=1}\right)}{\text{std}\left(\{R\left(\mathbf{x}^{i}_{T},c\right)\}^{N}_{i=1}\right)},$ (9)
It is important to note that, we have dropped the KL Loss. Although KL Loss can mitigate reward hacking to some extent (Liu et al., 2025a), inspired by yifan123 (2025), we use hybrid inference at test time, which can significantly address the reward hacking issue (See Appendix B).
MixGRPO reduces NFE of $\pi_{\theta}$ compared to all-timesteps optimization. However, the NFE of $\pi_{\theta_{\text{old}}}$ is not reduced, as complete inference is required to obtain the final image for reward calculation. In Section 3.3, we will introduce the use of higher-order ODE solvers, which also reduce the NFE of $\pi_{\theta_{\text{old}}}$ leading to further speedup. In summary, the mixed ODE-SDE sampling significantly reduces overhead while ensuring the introduction of randomness, allowing for optimization by GRPO.
@@ 3.2 Sliding Window as Optimization Scheduler
In fact, the interval $S$ can be non-fixed during the training process. In this section, we will introduce the sliding window to describe the movement of $S$, which leads to a significant improvement in the quality of the generated images. Along the denosing time-steps $\{0,1,\ldots,T-1\}$, MixGRPO defines a SDE sliding window $W(l)$ and optimization is only employed at the steps within $W(l)$.
@@
The left boundary $l$ of the sliding window moves as the training progresses. In the experiments, we found that the window size $w$, shift interval $\tau$, and window stride $s$ are all crucial hyperparameters. Through ablation studies (See Experiment 4.4.1), we identified the optimal settings. When the total sampling steps $T=25$, the best performance is achieved with $w=4$, $\tau=25$ and $s=1$. The detailed sliding window strategy and MixGRPO algorithm can be found in Algorithm 1.
Restricting the use of SDE sampling within the sliding window not only ensures the diversity of the generated images but also allows the model to concentrate on optimizing the flow within that window. Moving along the denoising direction represents the stochasticity of the probability flow from strong to weak, as illustrated in Figure 2. This is essentially a greedy strategy and is similar to the RL that assigns discount factors to process rewards (Pitis, 2019; Amit et al., 2020; Hu et al., 2022b), which gives greater importance to rewards derived from a larger search space in the earlier process.
The experimental results of different movement strategies in Table 5 demonstrate the validity of this intuition. Not only does the progressive strategy outperform random selection, but we also found that even when the sliding window is not moved (frozen), meaning only the earlier timesteps are optimized, MixGRPO can still yield good results, particularly in terms of ImageReward (Xu et al., 2023) and UnifiedReward (Wang et al., 2025). Based on this intuition, we also proposed an exponential decay strategy as follows, allowing $\tau$ to decrease as the window moves, enabling the model to avoid excessive optimization in smaller search spaces.
@@
where $\tau_{0}$ is the initial shift interval, $k$ is the decay factor, and $\lambda_{thr}$ is the threshold that controls when the decay starts. The exponential function $\exp(x)$ calculates $e^{x}$, while the Rectified Linear Unit $\text{ReLU}(x)$ is defined as $\max(0,x)$. Table 5 shows that the exponential decay strategy can achieve better results in terms of Pick Score (Kirstain et al., 2023) and ImageReward (Xu et al., 2023). This may be because the model focuses on the earlier stages of denoising, which can lead to more significant high-level changes, precisely what the human preference alignment reward model emphasizes.
@@ 3.3 Trade-off Between Overhead and Performance
MixGRPO employs SDE sampling within the sliding window and ODE sampling outside of it, allowing the use of higher-order ODE solvers to accelerate GRPO training-time sampling. The timesteps that utilize ODE sampling are divided into those before and after the sliding window. The timesteps after the sliding window solely influence the reward calculation, whereas the timesteps before the window affect both the reward and contribute to cumulative errors in the policy ratio computing. Therefore, we focus exclusively on the acceleration of the timesteps after the window.
Gao et al. (2024) has demonstrated the equivalence between the ODE sampling of flow matching models (FM) and DDIM, and Section 3.1 has also shown that diffusion probabilistic models (DPM) and FM share the same ODE form during the denoising process. Therefore, the higher-order ODE solvers e.g., DPM-Solver Series (Lu et al., 2022a; b; Zheng et al., 2023), UniPC (Zhao et al., 2023) designed for DPM sampling acceleration are also applicable to FM. We have reformulated DPM-Solver++ (Lu et al., 2022b) to apply it in the FM framework for ODE sampling acceleration and released detailed derivations in Appendix C.
By applying higher-order solvers, we achieve acceleration in the sampling of $\pi_{\theta_{\text{old}}}$ during GRPO training, which is essentially a balance between overhead and performance. Excessive acceleration leads to fewer timesteps, which inevitably results in a decline in image generation quality, thereby accumulating errors in the computation of rewards. We have found in practice that the 2nd-order DPM-Solver++ is sufficient to provide significant acceleration while ensuring that the generated images align well with human preferences in Table 9. Ultimately, we introduced DPM-Solver++ and adopted both progressive and frozen sliding window strategies, proposing MixGRPO-Flash and MixGRPO-Flash*. A detailed description of the algorithm can be found in Appendix E. These approaches achieve a greater degree of acceleration compared to MixGRPO, while also outperforming DanceGRPO in terms of human preference alignment performance.
@@ 4.1 Experiment Setup
Dataset  We conduct experiments using the prompts provided by the HPDv2111https://huggingface.co/datasets/ymhao/HPDv2 dataset, which is the official dataset for the HPS-v2 benchmark (Wu et al., 2023). The training set contains 103,700 prompts; in fact, MixGRPO achieved good human preference alignment results after training one epoch with only 9,600 prompts. The test set consists of 400 prompts. The prompts are diverse, encompassing four styles: “Animation”, “Concept Art”, “Painting”, and “Photo”.
Model  Following DanceGRPO (Xue et al., 2025), we use FLUX.1 Dev (Labs, 2024) as the base model, which is an advanced text-to-image model based on flow matching.
Overhead Evaluation  For the evaluation of overhead, we use two metrics: the number of function evaluations (NFE) (Lu et al., 2022a) and the time consumption per iteration during training. The NFE is divided into $\text{NFE}_{\pi_{\theta_{\text{old}}}}$ and $\text{NFE}_{\pi_{\theta}}$. $\text{NFE}_{\pi_{\theta_{\text{old}}}}$ represents the number of forward propagation of the reference model for computing the policy ratio and generating images. $\text{NFE}_{\pi_{\theta}}$ is the number of forward propagation of the policy model solely for the policy ratio. Additionally, the average training time per GRPO iteration provides a more accurate reflection of the acceleration effect.
Performance Evaluation  We used four multiple reward models in conjunction for GRPO, namely HPS-v2.1 (Wu et al., 2023), Pick Score (Kirstain et al., 2023), ImageReward (Xu et al., 2023) and Unified Reward Wang et al. (2025), both as reward guidance during training and as evaluation metrics. These metrics are all based on human preferences but emphasize different aspects. For example, ImageReward (Xu et al., 2023) highlights image-text alignment and fidelity, while Unified Reward (Wang et al., 2025) concentrates on semantics. DanceGRPO Xue et al. (2025) also demonstrates that using multiple reward models can achieve better results. To validate the robustness of MixGRPO, we also followed DanceGRPO and conducted additional comparisons using HPS-v2.1 as a single reward, and combining HPS-v2.1 (Wu et al., 2023) and CLIP Score (Radford et al., 2021) as multi-rewards.
@@ 4.2 Implementation Details
For training, all experiments are conducted on 32 Nvidia GPUs with a batch size of 1 and a maximum of 300 iterations. We use AdamW (Loshchilov & Hutter, 2017) as the optimizer with a learning rate of 1e-5 and a weight decay coefficient of 0.0001. Mixed precision training is used with the bfloat16 (bf16) format, while the master weights are maintained at full precision (fp32).
4.3 Main Experiments
In the main experiment, the four human-preference-based rewards were aggregated according to advantages, as shown in Algorithm 1. We evaluated the overhead and performance of MixGRPO in comparison with DanceGRPO, with the results presented in Table 1. The official DanceGRPO uses $\text{NFE}_{\pi_{\theta}}=14$; however, for fairness, we also tested DanceGRPO with $\text{NFE}_{\pi_{\theta}}=4$. For MixGRPO-Flash, we evaluated both the progressive and frozen strategies, and to ensure fairness, we also applied the frozen strategy to DanceGRPO.
We selected multiple scene prompts to visualize the results for FLUX.1 Dev, the officially configured DanceGRPO, and MixGRPO, as shown in Figure 3. It can be observed that MixGRPO achieved the best results in terms of semantics, aesthetics, and text-image alignment. Figure 5 shows the comparison results of DanceGRPO, MixGRPO, and MixGRPO-Flash with $\text{NFE}_{\pi_{\theta}}=4$. It can be observed that under the same overhead conditions, MixGRPO achieved better results compared to DanceGRPO. Additionally, MixGRPO-Flash enables accelerated sampling of $\pi_{\theta_{\text{old}}}$, and as the overhead decreases, the quality of the generated images still maintains a strong alignment with human preferences.
Following DanceGRPO (Xue et al., 2025), we also trained and evaluated the model using a single reward model, e.g., HPSv2.1 and two reward models, e.g., HPSv2.1 and CLIP Score on the HPDv2 dataset (Wu et al., 2023). The results (See Table 3) demonstrate that MixGRPO achieves the best performance on both in-domain and out-of-domain rewards, whether using a single or multiple reward models. The visualized results are displayed in Figure 6 and Figure 7 of Appendix F.
To demonstrate the robustness of the MixGRPO, we followed Flow-GRPO (Liu et al., 2025a) and applied the Low-Rank Adaptation (LoRA) (Hu et al., 2022a) method on the Stable Diffusion 3.5 (SD3.5) (Esser et al., 2024). We used HPS-v2.1 (Wu et al., 2023), Pick Score (Kirstain et al., 2023), and ImageReward (Xu et al., 2023) as multi-rewards, comparing Flow-DPO (offline/online) (Liu et al., 2025b), and Flow-GRPO (Liu et al., 2025a). The results in Table 3 showed that MixGRPO is well-suited for the LoRA, significantly reducing training overhead while outperforming both Flow-DPO and Flow-GRPO in terms of speed and performance. The visualization results can be found in Figure 8 of Appendix F.
Table 2: Comparison results demonstrate that MixGRPO achieves the best performance on both in-domain and out-of-domain rewards.
@@ 4.3 Main Experiments
Table 3: Comparison results demonstrate that MixGRPO outperforms Flow-DPO and Flow-GRPO in training efficiency and performance.
@@ 4.4.1 Sliding Window Hyperparamters
As introduced in Section 3.2, the moving strategy, shift interval $\tau$, window size $w$ and window stride $s$ are all important parameters of the sliding window. We conducted ablation experiments on each of them. For the moving strategy, we compared three approaches: frozen, where the window remains stationary; random, where a random window position is selected at each iteration; and progressive, where the sliding window moves incrementally with the denoising steps. For the progressive strategy, we tested different scheduling strategies where the interval $\tau$ initially starts at 25 but changes with training iterations. As shown in Table 5, the results indicate that under the progressive strategy, either exponential decay or constant scheduling strategies are optimal. For the shift interval $\tau$, 25 is the optimal setting (See Table 5).
The number of inferences for $\pi_{\theta}$ increases with the growth of the window size $w$, leading to greater time overhead. We compared different settings of $w$, and the results are shown in Table 7. Ultimately, we selected $w=4$ as a balanced setting between overhead and performance. For the window stride $s$, we found through experimentation that $s=1$ is the optimal choice, as shown in Table 7.
@@
Strategy
Interval
Schedule
HPS-v2.1 Pick Score ImageReward Unified Reward
@@
Table 5: Comparison for shift interval $\tau$.
@@
Table 6: Comparison for window size $w$
@@
Table 7: Comparison for window stride $s$
@@
4.4.2 High Order ODE Solver
MixGRPO enables the possibility of accelerating ODE sampling with high-order ODE solvers by combining SDE and ODE sampling methods. We first conducted ablation experiments on the order of the solver, using DPM-Solver++ (Lu et al., 2022b) as the high-order solver with the progressive strategy. The results, as shown in Table 9, indicate that the second-order mid-point method is the optimal setting. Then, as described in Section 3.3, we compared two acceleration approaches. One is MixGRPO-Flash, which utilizes the progressive window moving strategy. The other is MixGRPO-Flash*, which employs the frozen moving strategy. They all achieve a balance between overhead and performance by reducing the number of ODE sampling steps after the sliding window. However, in practice, MixGRPO-Flash requires the window to move throughout the training process, which results in a shorter ODE portion being accelerated. Consequently, the acceleration effect of MixGRPO-Flash, on average, is not as pronounced as that of MixGRPO-Flash*.
Table 8: Comparison of performance across different-order solvers. The second-order Midpoint method achieves the best performance.
@@ 4.4.2 High Order ODE Solver
Table 9: Comparison of different sampling steps for $\pi_{\theta_{\text{old}}}$. MixGRPO-Flash* achieves good performance even with few steps.

### v3

@@ 3.1 Mixed ODE-SDE Sampling in GRPO
MixGRPO reduces the NFE of $\pi_{\theta}$ and optimized timesteps count compared to global SDE sampling. However, the NFE of $\pi_{\theta_{\text{old}}}$ is not reduced, as complete inference is required to obtain the final image for reward calculation. In Section 3.3, we will introduce the use of higher-order ODE solvers, which also reduce the NFE of $\pi_{\theta_{\text{old}}}$ leading to further speedup. In summary, the mixed ODE-SDE sampling significantly streamlines the MDP, enhancing training efficiency by lowering optimization overhead and enabling more focused gradient updates.
Algorithm 1 MixGRPO Training Process
3:Init left boundary of $W(l)$: $l\leftarrow 0$
4:for training iteration $m=1$ to $M$ do
5:  Sample batch prompts $\mathcal{C}_{b}\sim\mathcal{C}$
7:  for each prompt $\mathbf{c}\in\mathcal{C}_{b}$ do
8:    Init the same noise $\mathbf{x_{0}}\sim\mathcal{N}(0,\mathbf{I})$
9:    for generate $i$-th image from $i=1$ to N do
10:     for sampling timestep $t=0$ to T1T-1 do
11:      if $t\in W(l)$ then
12:         Use SDE Sampling to get $\mathbf{x}^{i}_{t+1}$
13:      else
14:         Use ODE Sampling to get $\mathbf{x}^{i}_{t+1}$
15:      end if
16:     end for
17:    end for
18:    for $i$-th image from $i=1$ to N do
19:     Calculate advantages: $A_{i}\leftarrow\sum_{k=1}^{K}\frac{R(\mathbf{x}^{i}_{T},\mathbf{c})^{i}_{k}-\mu_{k}}{\sigma_{k}}$
20:    end for
21:    for optimization timestep $t\in W(l)$ do
22:     Update policy model: $\theta\leftarrow\theta+\eta\nabla_{\theta}\mathcal{J}$
23:    end for
24:  end for
26:    $l\leftarrow\min(l+s,~T-w)$
27:  end if
28:end for
@@ 3.2 Sliding Window as Optimization Scheduler
In this section, we will introduce the sliding window to describe the movement of $S$, which leads to a significant improvement in the quality of the generated images. Along the denosing time-steps $\{0,1,\ldots,T-1\}$, MixGRPO defines a SDE sliding window $W(l)$ and optimization is only employed at the timesteps within $W(l)$.
@@
The left boundary $l$ of the sliding window moves as the training progresses. The movement strategy of the sliding window is governed by the following parameters. $\tau$ is the shift interval, which specifies the number of training steps between each window shift. $s$ is the window stride, which indicates the number of denoising timesteps that the left boundary $l$ advances during each shift. The detailed sliding window strategy and MixGRPO algorithm can be found in Algorithm 1. In Section 4.4.1, we conduct thorough ablation studies to validate these parameters.
In our MixGRPO, the denoising process inherently follows a stochastic-to-deterministic trajectory: moving along the probability flow corresponds to a transition from high to low stochasticity, as illustrated in Figure 2 (right). Our sliding-window approach operationalizes this principle by creating an implicit training curriculum. This is analogous to reward shaping in RL [26, 2, 10], where discount factors prioritize rewards from the most critical steps in a process. For diffusion and flow-matching models, this curriculum directs the GRPO algorithm to first optimize low-SNR regions—where dynamics are highly stochastic—before shifting focus to the more deterministic, high-SNR regions for fine-grained refinement. This coarse-to-fine strategy stabilizes the initial, more volatile stages of training and enables the model to learn a robust global structure before perfecting finer details.
The experimental results of different movement strategies in Table 4 demonstrate the validity of this intuition. Not only does the progressive strategy outperform random selection, but our findings also indicate that MixGRPO can achieve strong performance even with a static (frozen) window. In this setting, where optimization is exclusively applied to the initial timesteps, our method still surpasses DanceGRPO. This is because the initial timesteps in the generative process, which restores an image from Gaussian noise, account for a more substantial degree of denoising than other timesteps [11, 4]. Initial timesteps introduces greater stochasticity into the final output, thereby increasing the likelihood of generating images that align with the preferences of a reward model and ultimately yielding a higher optimization payoff. Based on this intuition, we also proposed an exponential decay strategy as follows, allowing $\tau$ to decrease as the window moves, enabling the model to avoid excessive optimization in smaller search spaces.
@@
where $\tau_{0}$ is the initial shift interval, $k$ is the decay factor, and $\lambda_{thr}$ is the threshold that controls when the decay starts. The exponential function $\exp(x)$ calculates $e^{x}$, while the Rectified Linear Unit $\text{ReLU}(x)$ is defined as $\max(0,x)$. Table 4 shows that the exponential decay strategy can achieve better results in terms of Pick Score [12] and ImageReward [41]. This may be because the model focuses on the earlier stages of denoising, which can lead to more significant high-level changes, precisely what the human preference alignment reward model emphasizes.
@@ 3.3 Trade-off Between Overhead and Performance
Unlike DanceGRPO [42], which relies on global SDE sampling, MixGRPO employs a mixed ODE-SDE method, allowing the use of higher-order ODE solvers to accelerate GRPO training-time sampling. However, accelerating the ODE phase before the SDE window amplifies the solver’s numerical errors due to the window’s stochasticity, severely degrading image quality and corrupting the reward signal. In contrast, accelerating the ODE phase after the window provides a significant speed-up without compromising the image fidelity required for reliable reward evaluation. Therefore, MixGRPO exclusively accelerates the post-window ODE timesteps.
[7] has demonstrated the equivalence between the ODE sampling of flow matching models (FM) and DDIM, and Section 3.1 has also shown that diffusion probabilistic models (DPM) and FM share the same ODE form during the denoising process. Therefore, the higher-order ODE solvers e.g., DPM-Solver Series [21, 22, 47], UniPC [46] designed for DPM sampling acceleration are also applicable to FM. We have reformulated DPM-Solver++ [22] to apply it in the FM framework for ODE sampling acceleration and released detailed derivations in Supplementary.
By applying higher-order solvers, we achieve acceleration in the sampling of $\pi_{\theta_{\text{old}}}$ during GRPO training, which is essentially a balance between overhead and performance. Excessive acceleration leads to fewer timesteps, which inevitably results in a decline in image generation quality, thereby accumulating errors in the computation of rewards. We have found in practice that the 2nd-order DPM-Solver++ is sufficient to provide significant acceleration while ensuring that the generated images align well with human preferences in Table 8.
Ultimately, we introduced DPM-Solver++ and adopted both progressive and frozen sliding window strategies, proposing MixGRPO-Flash and MixGRPO-Flash*. A detailed description of the algorithm can be found in Supplementary. These faster versions offer more substantial acceleration relative to base version of our MixGRPO, while simultaneously surpassing DanceGRPO in their alignment with human preferences.
@@ 4.1 Experiment Setup
Dataset  We conduct experiments using the prompts provided by the HPDv2 dataset, which is the official dataset for the HPS-v2 benchmark [40]. The training set contains 103,700 prompts; in fact, MixGRPO achieved good human preference alignment results after training one epoch with only 9,600 prompts. The test set consists of 400 prompts. The prompts are diverse, encompassing four styles: “Animation”, “Concept Art”, “Painting”, and “Photo”.
Model  Following DanceGRPO [42], we use FLUX.1 Dev [13] as the base model, which is an advanced text-to-image model based on flow matching.
Overhead Evaluation  For the evaluation of overhead, we use two metrics: the number of function evaluations (NFE) [21] and the time consumption per iteration during training. The NFE is divided into $\text{NFE}_{\pi_{\theta_{\text{old}}}}$ and $\text{NFE}_{\pi_{\theta}}$. $\text{NFE}_{\pi_{\theta_{\text{old}}}}$ represents the number of forward propagation of the reference model for computing the policy ratio and generating images. $\text{NFE}_{\pi_{\theta}}$ is the number of forward propagation of the policy model solely for the policy ratio. Additionally, the average training time per GRPO iteration provides a more accurate reflection of the acceleration effect.
Performance Evaluation  We used four multiple reward models in conjunction for GRPO, namely HPS-v2.1 [40], Pick Score [12], ImageReward [41] and Unified Reward [37], both as reward guidance during training and as evaluation metrics. These metrics are all based on human preferences but emphasize different aspects. For example, ImageReward [41] highlights image-text alignment and fidelity, while Unified Reward [37] concentrates on semantics. DanceGRPO [42] also demonstrates that using multiple reward models can achieve better results. To validate the robustness of MixGRPO, we also followed DanceGRPO and conducted additional comparisons using HPS-v2.1 as a single reward, and combining HPS-v2.1 [40] and CLIP Score [27] as multi-rewards.
@@ 4.2 Implementation Details
For training, all experiments are conducted on 32 Nvidia GPUs with a batch size per gpu of 1 and a maximum of 300 iterations. We use AdamW [20] as the optimizer with a learning rate of 1e-5 and a weight decay coefficient of 0.0001. Mixed precision training is used with the bfloat16 (bf16) format, while the master weights are maintained at fp32 precision.
4.3 Comparison with existing methods
To validate the generality, effectiveness, and efficiency of our proposed method, we conducted comprehensive experiments across two distinct frameworks: Dance-GRPO and Flow-GRPO. In our experiments, we employed multiple reward models for guidance, as described in Algorithm 1.
We first evaluated the overhead and performance of MixGRPO in comparison with DanceGRPO, with the results presented in Table 1. The official DanceGRPO uses $\text{NFE}_{\pi_{\theta}}=14$; For fairness, we also tested DanceGRPO with $\text{NFE}_{\pi_{\theta}}=4$. For MixGRPO-Flash, we evaluated both the progressive and frozen strategies, and to ensure fairness, we also applied the frozen strategy to DanceGRPO. The visual results are presented in Figure 3. We observe that MixGRPO achieves the best performance in terms of semantics, aesthetics, and text-image alignment. For a direct comparison at $\text{NFE}_{\pi_{\theta}}=4$, the qualitative results are shown in Figure 4. confirms the robustness of acceleration in MixGRPO-Flash, demonstrating the model’s ability to preserve the aesthetic quality and semantic alignment while speeding up the process. Following DanceGRPO [42], we also trained and evaluated the model using a single reward model, e.g., HPSv2.1 and two reward models, e.g., HPSv2.1 and CLIP Score on the HPDv2 dataset [40]. The results (See Table 2) demonstrate that MixGRPO achieves the best performance on both in-domain and out-of-domain rewards, whether using a single or multiple reward models. The visualized results are displayed in Supplementary.
Table 2: Comparison with DanceGRPO. The results demonstrate that MixGRPO achieves the best performance on both in-domain and out-of-domain rewards.
@@ 4.3 Comparison with existing methods
Table 3: Comparison with FlowDPO and FLow-GRPO. The results demonstrate that MixGRPO outperforms others in training efficiency and performance.
@@
Methodologically, Flow-GRPO [17] and DanceGRPO [42] are conceptually similar, differing only in minor aspects such as their sampling formulations and the use of KL-regularization. Therefore, to ensure a fair and direct comparison, we benchmarked our method against their respective officially released models. To demonstrate the robustness of the MixGRPO across distinct foundational models, we also followed Flow-GRPO [17] and applied the Low-Rank Adaptation (LoRA) [9] on the Stable Diffusion 3.5(SD3.5) [4]. We used HPS-v2.1 [40], Pick Score [12], and ImageReward [41] as multi-rewards, comparing Flow-DPO (offline/online) [18], and Flow-GRPO [17]. The results in Table 3 showed that MixGRPO is well-suited for the LoRA, significantly reducing training overhead while outperforming both Flow-DPO and Flow-GRPO in terms of speed and performance. The visualization results can be found in Supplementary.
@@ 4.4.1 Sliding Window Hyperparamters
Strategy Interval Schedule HPS-v2.1 Pick Score ImageReward Unified Reward
@@
Table 5: Comparison for different shift interval $\tau$.
@@
As introduced in Section 3.2, the moving strategy, shift interval $\tau$, window size $w$ and window stride $s$ are parameters introduced by the SDE sliding window. We conducted ablation experiments on each of them.
For the moving strategy, we compared three approaches: frozen, where the window remains stationary; random, where a random window position is selected at each iteration; and progressive, where the sliding window moves incrementally with the denoising steps. For the progressive strategy, we tested different scheduling strategies where the interval $\tau$ initially starts at 25 but changes with training iterations. As shown in Table 4, the results indicate that under the progressive strategy, either exponential decay or constant scheduling strategies are optimal. For the shift interval $\tau$, 25 is the optimal setting (See Table 5). The number of inferences for $\pi_{\theta}$ increases with the growth of the window size $w$, leading to greater time overhead. We compared different settings of $w$, and the results are shown in Table 6. Ultimately, we selected $w=4$ as a balanced setting between overhead and performance. For the window stride $s$, we found through experimentation that $s=1$ is the optimal choice, as shown in Table 7.
In practice, the optimal settings are applicable not only to Flux [13] but also to Stable Diffusion 3.5 [4]. Moreover, the model can maintain its optimal generation performance within a controllable range of parameter adjustments. This demonstrates the robustness of the sliding-window optimization scheduling strategy.
Table 6: Comparison for different window size $w$.
@@
Table 7: Comparison for different window stride $s$.
@@
Table 8: Comparison of performance across different solvers. The second-order Midpoint method achieves the best performance.
@@
4.4.2 High Order ODE Solver
MixGRPO-Flash, which incorporates a high-order ODE solver to accelerate the sampling process of the reference model, achieves an effective trade-off between speed and performance. For MixGRPO-Flash, we first conducted ablation experiments on the order of the solver, using DPM-Solver++ [22] as the high-order solver with the progressive strategy. The results, as shown in Table 8, indicate that the second-order mid-point method is the optimal setting, optimizing the most human preference alignment metrics while simultaneously accelerating the process.
Then we compared two acceleration approaches. One is MixGRPO-Flash, and the other is MixGRPO-Flash*. Both utilize a second-order ODE solver for acceleration, but they differ in their sliding-window moving strategies. The quantitative results are presented in Table 9. MixGRPO-Flash requires the window to move throughout the training process, resulting in a smaller portion of the ODE being accelerated compared to MixGRPO-Flash*. Consequently, MixGRPO-Flash* not only achieves a higher degree of acceleration for the reference model but also yields superior results in the ImageReward [41] and Unified Reward [37] metrics.
Figure 4: Qualitative results of MixGRPO-Flash. As the training overhead is reduced, generated images maintain high quality.
Table 9: Comparison about different sampling steps of the reference model. MixGRPO-Flash and MixGRPO-Flash* all achieve a balance between overhead and performance.